# Optimal In-Network Distribution of Learning Functions for a Secure-by-Design Programmable Data Plane of Next-Generation Networks

### v1

@@ Abstract
The rise of programmable data plane (PDP) and in-network computing (INC) paradigms paves the way for the development of network devices (switches, network interface cards, etc.) capable of performing advanced computing tasks. This allows to execute algorithms of various nature, including machine learning ones, within the network itself to support user and network services. In particular, this paper delves into the issue of implementing in-network learning models to support distributed intrusion detection systems (IDS). It proposes a model that optimally distributes the IDS workload, resulting from the subdivision of a “Strong Learner” (SL) model into lighter distributed “Weak Learner” (WL) models, among data plane devices; the objective is to ensure complete network security without excessively burdening their normal operations. Furthermore, a meta-heuristic approach is proposed to reduce the long computational time required by the exact solution provided by the mathematical model, and its performance is evaluated. The analysis conducted and the results obtained demonstrate the enormous potential of the proposed new approach to the creation of intelligent data planes that effectively act as a first line of defense against cyber attacks, with minimal additional workload on network devices.
@@
In-Network Computing, Distributed AI, IDS, Programmable Data Plane, Security by Design.
@@ I Introduction
The evolving cyber threat landscape requires increasingly agile and adaptable cyber-security solutions. The emerging paradigms of in-network computing (INC) and in-network distributed learning (INDS), coupled with the concept of distributed Intrusion Detection Systems (IDS), emerge as key components to address the challenge. The integration of these concepts has in fact the potential to revolutionize network security by offering a robust, scalable, and resilient defense against ever-evolving threats.
INC exploits the idea of distributing computational tasks across the network infrastructure, rather than relying solely on edge or cloud computing resources. To this end, it leverages the capabilities of network devices, such as switches, routers, and network interface cards (NICs), to perform data processing or caching. An interesting subfield of In-Network Computing that focuses on the use of distributed artificial intelligence (AI) techniques is the so-called “In-network distributed intelligence”, which aims to enable network devices to collaborate and make intelligent decisions autonomously, without the need for centralized control. This paradigm can make networks more scalable and fault-tolerant (as they become less dependent on centralized controls) and highly adaptable to changing conditions and traffic distributions in real-time through intelligent decisions about traffic routing, resource management, and network performance optimization.
@@
For their part, the fixed-perimeter nature of traditional IDSs is no longer adequate for the highly pervasive and dynamic nature of next-generation networks. Even recent solutions in the literature, which rely on in-network telemetry and traffic data forwarding to a centralized SDN controller that runs the detection module and completes the decision-making process, do not meet the mentioned requirements.
Next-generation networks require Active IDSs (also called Intrusion Prevention Systems - IPS), which leverage the INC and distributed intelligence paradigms to process and analyze network data within Programmable Data Plane (PDP) devices, and enable the devices themselves to block threats through completely decentralized procedures; thereby improving the effectiveness and timeliness of intrusion detection and ensuring greater scalability, resilience, and fault tolerance.
@@
For the aforementioned paradigm to be truly effective, orchestration is required to always implement an optimal distribution of learning functions that truly allows the network to (i) continuously improve the accuracy of intrusion detection by adapting to new threats, (ii) reduce the processing load, and (iii) reduce both the impact on the standard functionality of the involved network devices (e.g., packet forwarding) and the reaction time to threats. The main contributions of this paper can therefore be summarized as follows:
  • demonstrate the potential of jointly using PDP devices and in-network distributed learning to enable the network user plane to implement a fully distributed active IDS, and increase the effectiveness of this new functionality;
  • propose an optimization model for efficient deployment of in-network learning models for distributed Active IDS, which balances security coverage with performance;
@@
  • conduct a comprehensive performance analysis aimed at demonstrating the effectiveness of the proposed approach in enhancing the protection of the network against cyber threats while minimizing the impact on the overall network performance.
The remainder of the paper is organized as follows. Section II presents the main related works in the key reference areas of this research. In Section III, an innovative paradigm that exploits distributed in-network learning models to implement a “secure-by-design” data plane is introduced, while Section IV illustrates a model for the optimization of the in-network distribution of learning elements and related meta-heuristic solution. The results of a comprehensive performance evaluation campaign are presented in Section V. Finally, in Section VI, conclusions are drawn and future work is outlined.
@@ II Related Works
II-A In-Network Security: ML/DL-aided Traffic Classification
With the advent of Programmable Data Plane (PDP) and INC capabilities, recent efforts have focused on the design of in-network IDS solutions (also referred to as in-network classifiers) to address security-related challenges. A significant area of research investigated the use of the programmable PISA (Protocol Independent Switch Architecture) switch architecture by means of Reconfigurable Match Tables (RMT), enabled by the introduction of the P4 language [2]. In [3] the authors proposed N2Net, a solution that implements the forwarding pass of a Binary Neural Network (BNN) in a P4-enabled switch, outlining the limitations of modern programmable networking devices in accommodating complex ML/DL models characterized by intricate computations and mathematical operations. Following this direction, the authors of BaNaNa Split [4] extended the use of the BNN to SmartNICs to overcome the mentioned limitations: the joint work of programmable networking devices and end-host applications. Nevertheless, the proposed solution does not fit well the concept of ubiquitous and pervasive in-network security, since it does not work without a server that shares the workload with the networking device.
@@ II-A In-Network Security: ML/DL-aided Traffic Classification
Parallel efforts have focused on encoding ML models within programmable networking devices, particularly Random Forests (RFs) and Decision Trees (DTs). In this direction, SwitchTree [7] and Forest [8] stand out as the most valuable examples. Both proposals strove to find the best encoding methodology to embed DTs and RFs within constrained and instruction set-limited PDPs. Following this trend, the works in [9, 10, 11, 12] show effort in designing a framework capable of encoding general RF/DT within P4-enabled networking devices. Recent research has demonstrated the remarkable capabilities of eBPF (extended Berkeley Packet Filter), showing nearly equivalent performance to P4 in managing general-purpose tasks offloaded to networking devices [13]. An important contribution in this domain is found in [14], where the authors focus on developing an efficient and effective encoding of a DNN using eBPF technology.
@@
A common effort emerging from the literature is the search for optimal encodings of the entire (sometimes complex) ML/DL models to adapt them to network devices with reduced impact on packet forwarding performance. None of them addresses how to intelligently distribute in-network classification modules to achieve pervasive and ubiquitous security through a fully distributed and collaborative approach of such modules, which is the objective of the novel paradigm studied in our paper.
II-B In-Network Learning Distribution
The paradigm of the distribution of computational functions relevant to AI (both training and inference) finds its first evidence in the context of Edge and Cloud Computing.
Many works in the literature addressed the concept of decomposing a deep neural network (DNN) into its layers to distribute the workload between an edge mobile device and the cloud, proposing optimization models for this purpose. Among others, in [15] the best split is determined via regression models that predict the computational and energy consumption of each DNN layer, while in [16] the optimal solution is determined by considering device and network resource utilization to minimize end-to-end latency between the edge and the cloud.
Only recently, with the emergence of the potential of the in-network computing paradigm [17], attention has shifted towards a distribution of learning functions that also exploits the network segments that connect Edge and Cloud. Understanding the close and crucial integration between artificial intelligence and future 6G networks, the authors of [18], [19] and [20] envisaged and analyzed the structural changes needed for the future 6G networks to naturally accommodate distributed artificial intelligence activities within their Data Plane.
Instead, Saquetti et. al. [21] focus on the constrained nature of PDP devices as well as the limitations imposed by the reference PDP programming language (i.e., P4) when dealing with distributed intelligence in the network. Through a simple PoC – a neural network with 3 layers and a total of seven neurons – they proposed an optimization model to distribute the DNN within the network at single neuron granularity, with a one-to-one mapping between PDP and neuron. However, it turns out that this type of distribution is not feasible when the neural network is complex, severely limiting the applicability of the proposal.
In the wake of the recent effort to deploy intelligence “in-the-network” by leveraging key enablers envisioned for upcoming 6G networks, our research aims to help fill a crucial literature and structural gap regarding network security for future generation networks. The close integration between AI and networks is a key factor for pursuing the concept of integrated security. By deploying virtualized anomaly detection and response functions across the network and enabling their collaborative action, a security fabric can be created that makes the network the first line of defense against malicious attempts. We think that this approach is essential to avoid the mistakes made with previous generations of networks, in which security was not designed in perfect synergy with the network itself but was treated as an “additional” functionality, thus opening the door to more intelligent and malicious attacks.
The potential of the approach described is accompanied by new challenges, such as finding the optimal positioning within the network of the AI-empowered security capabilities mentioned above to minimize both the delay in completing tasks and the resource consumption of the network devices involved. Our paper aims precisely to contribute to finding a solution to this compelling research problem.
III Deployment of an ML-enabled Active IDS in a network dataplane
The reference for the research reported in this paper is the one presented by the authors in [1], where a new paradigm according to which anomaly detection capabilities are natively embedded in the devices of a typical data plane of a future programmable network is introduced. That work in fact reports only a simple proof-of-concept of the resulting ML-enabled Active Intrusion Detection System, for which instead in the present paper we propose an effective method of optimizing the deployment of learning functions in the devices and their related chaining. For the benefit of the reader, we briefly report the basic concepts, referring to the aforementioned paper for the details of the hypothesized architecture.
@@ III-A Projecting the Ensemble Learning over the Network
The proposed paradigm envisages that through ad hoc functions included in the first level, the model that must be embedded in the PDP is trained, its partitioning is performed, and the VNFs that will carry out detection and response to attacks are created. An SL appropriately trained for the purpose is then broken down into individual WLs coded as WL-VNFs and made available to the orchestration functions that are in the second level. This process is depicted in Fig.1.
An optimal distribution strategy of the WL-VNFs among PDP devices is then decided, which allows the selected switches that host the WL-VNFs to operate cooperatively as an active IDS in the network. The activities described in this paper refer to what is only theorized at the second level of the mentioned architecture, but not previously developed. Specifically, the goal is to find the set of WL-VNFs and the switches that host them in such a way as to maximize the security coverage of the considered network, i.e., the effectiveness in detecting and reacting to the maximum number of attacks.
@@
The flow, as it passes through the network, is analyzed by the various WLs that constitute the SL, and each one signals the result of its inference to the others. If a WL realizes that it is the last of the set that constitutes an SL and that all the others have already performed the flow analysis, it completes its analysis, and through a majority voting algorithm takes the final decision, blocking the flows that the WLs chain deems malicious. The algorithm is completely distributed and does not require human involvement or of the network controller. To allow distributed WL-VNFs to inform each other on the inferences carried out for a network flow, a custom header, P4-encoded, is considered as well as a procedure carried out by the PDP device augmented with the WL-VNFs.
IV Formulation
@@
We represent a network using a graph. The nodes in our model represent the network nodes in which the WL-VNFs can be deployed, while the edges denote the links between network elements. We use node coloring to represent the implementation of specific WL-VNFs, where each color [^1] The terms “color” and “SL/WL-VNF” will be used interchangeably. More specifically, SL-VNF refers to a scenario in which only one color is needed. corresponds to a different type of WL-VNF and the coloring cost corresponds to the associated implementation cost. For instance, an SL composed of three WLs will determine three WL-VNFs and therefore three different colors (e.g., red, green, and blue), as shown in Fig.2.

### v2

@@ Abstract
The rise of programmable data plane (PDP) and in-network computing (INC) paradigms paves the way for the development of network devices (switches, network interface cards, etc.) capable of performing advanced processing tasks. This allows running various types of algorithms, including machine learning, within the network itself to support user and network services. In particular, this paper delves into the deployment of in-network learning models with the aim of implementing fully distributed intrusion detection systems (IDS) or intrusion prevention systems (IPS). Specifically, a model is proposed for the optimal distribution of the IDS/IPS workload among data plane devices with the aim of ensuring complete network security without excessively burdening the normal operations of the devices. Furthermore, a meta-heuristic approach is proposed to reduce the long computation time required by the exact solution provided by the mathematical model and its performance is evaluated. The analysis conducted and the results obtained demonstrate the enormous potential of the proposed new approach for the creation of intelligent data planes that act effectively and autonomously as the first line of defense against cyber attacks, with minimal additional workload on the network devices involved.
@@
In-Network Computing, Distributed AI, IDS, IPS Programmable Data Plane, Security by Design.
@@ I Introduction
The evolving cyber threat landscape requires increasingly agile and adaptable cyber-security solutions.
In this context, a suitable integration of emerging paradigms, such as In-Network Computing (INC) and Distributed Intelligence, has the potential to revolutionize network security by offering robust, scalable, and resilient defense solutions against ever-evolving threats.
In-Network Computing exploits the idea of distributing computational tasks across the network infrastructure, rather than relying solely on edge or cloud computing resources. To this end, it leverages the capabilities of data plane network devices, such as switches, routers, and network interface cards (NICs), to perform data processing or caching. An interesting subfield of In-Network Computing that focuses on the use of distributed artificial intelligence (AI) techniques is the so-called “In-network distributed intelligence”, which aims to enable network devices to collaborate and make intelligent decisions autonomously, without the need for centralized control. This paradigm can make networks more scalable and fault-tolerant (as they become less dependent on centralized controls) and highly adaptable to changing conditions and traffic distributions in real-time through intelligent decisions about traffic routing, resource management, and network performance optimization.
@@
As far as network security issues are concerned, the fixed-perimeter nature of traditional IDS and IPS solutions is no longer adequate for the highly pervasive and dynamic nature of next-generation networks. Perimeter IDS and IPS in fact fail to be effective in case of attacks coming from inside the network that are continuously increasing. This is a consequence of the expansion of the attack surface resulting from the growing number of devices that will access next-generation networks in the future (IoT devices, vehicles, household appliances, etc.)
Even recent solutions, which rely on in-network telemetry and traffic data forwarding to a centralized SDN controller that runs the detection module and completes the decision-making process are not adequate to future networking scenarios. These solutions may leverage sophisticated AI mechanisms to process and analyze network data in a centralized manner, with obvious inefficiencies in terms of increased control traffic and poor scalability. Alternative mechanisms leveraging distributed intelligence to process and analyze network data within Programmable Data Plane (PDP) devices, and enable the devices themselves to block threats through completely decentralized procedures can definitely improve the effectiveness and timeliness of intrusion detection, while ensuring greater scalability, resilience, and fault tolerance.
@@
For the aforementioned paradigm to be truly effective, orchestration is required to always implement an optimal distribution of learning functions that truly allows the network to experience (i) a good level of threat protection , (ii) a reduced processing load, and (iii) a reduced impact on the standard functionality of the involved network devices (e.g., packet forwarding) and a reduced reaction time to threats.
The main contributions of this paper can therefore be summarized as follows:
  • propose an optimization model for efficient deployment of in-network learning models for distributed IPS, which balances security coverage with performance;
@@
  • conduct a comprehensive performance analysis aimed at demonstrating the effectiveness of the proposed approach in enhancing the protection of the network against DDoS cyber threats while minimizing the impact on the overall network performance.
The remainder of the paper is organized as follows. Section II presents the main related works in the key reference areas of this research. In Section III, the reference paradigm that exploits distributed in-network learning models to implement a “secure-by-design” data plane is introduced, while Section IV illustrates a model for the optimization of the in-network distribution of learning elements and related meta-heuristic solution. The results of a comprehensive performance evaluation campaign are presented in Section V. Finally, in Section VI, conclusions are drawn and future work is outlined.
@@ II Related Works
Recent works in the literature address the issue of distributing computational functions relevant to AI (both training and inference), aiming to decompose a Deep Neural Network (DNN) into its layers to distribute the workload among a mobile Edge device and the Cloud. Among the optimization models proposed for this purpose, in [2] the best split is determined by regression models predicting the computational and energy consumption of each DNN layer, while in [3] the optimal solution is determined considering the network and device resources usage to minimize the end-to-end latency between edge and cloud.
With the emergence of the potential of the in-network computing paradigm [4], the focus has shifted towards a distribution of learning functions that also leverages the network segments connecting Edge and Cloud, namely Programmable Data Plane (PDP) and INC capabilities. Realizing the tight and crucial integration between AI and future 6G networks, the authors of [5], [6] and [7] have envisioned and analyzed the structural changes needed for future 6G networks to naturally accommodate distributed AI tasks within their Data Plane.
A significant area of research investigated the use of the programmable PISA (Protocol Independent Switch Architecture) switch architecture by means of Reconfigurable Match Tables (RMT), enabled by the introduction of the P4 language [8]. In [9] the authors proposed N2Net, a solution that implements the forwarding pass of a Binary Neural Network (BNN) in a P4-enabled switch, outlining the limitations of modern programmable networking devices in accommodating complex ML/DL models characterized by intricate computations and mathematical operations. Following this direction, the authors of BaNaNa Split [10] extended the use of the BNN to SmartNICs to overcome the mentioned limitations: the joint work of programmable networking devices and end-host applications. Nevertheless, the proposed solution does not fit well the concept of ubiquitous and pervasive in-network security, since it does not work without a server that shares the workload with the networking device. Instead, Saquetti et. al. [11] focus on the constrained nature of PDP devices as well as the limitations imposed by the reference PDP programming language (i.e., P4) when dealing with distributed intelligence in the network. Through a simple PoC – a neural network with 3 layers and a total of seven neurons – they proposed an optimization model to distribute the DNN within the network at single neuron granularity, with a one-to-one mapping between PDP and neuron. However, it turns out that this type of distribution is unfeasible when the NN is complex, severely limiting the applicability of the proposal.
@@
Parallel efforts have focused on encoding ML models within programmable networking devices, particularly Random Forests (RFs) and Decision Trees (DTs). In this direction, SwitchTree [14] and Forest [15] stand out as the most valuable examples. Both proposals strove to find the best encoding methodology to embed DTs and RFs within constrained and instruction set-limited PDPs. Following this trend, the works in [16, 17, 18, 19] show effort in designing a framework capable of encoding general RF/DT within P4-enabled networking devices.
@@
A common effort emerging from the literature is the search for optimal encodings of the entire (sometimes complex) ML/DL models to adapt them to network devices with reduced impact on packet forwarding performance. None of them addresses how to intelligently distribute in-network classification modules to achieve pervasive and ubiquitous security through a fully distributed and collaborative approach of such modules, which is the objective of the novel paradigm studied in our paper. The potential of the approach described is accompanied by new challenges, such as finding the optimal positioning within the network of the AI-empowered security capabilities mentioned above to minimize both the delay in completing tasks and the resource consumption of the network devices involved. Our paper aims precisely to contribute to finding a solution to this compelling research problem.
III Deployment of an ML-enabled IPS in a network dataplane
The reference framework for the research reported in this paper is the one presented by the authors in [20], where a new paradigm, according to which anomaly detection capabilities are deployed in a fully distributed way within devices of a typical data plane of a future programmable network, is introduced. Only the proposed innovative architecture and a simple proof-of-concept study were the subject of that paper in which among the most important future issues highlighted and left open there is the design of an optimization policy for the distribution of the learning functions within the most suitable data plane devices and their related chaining. The latter key issue is precisely the one addressed in the present paper.
@@ III-A Projecting the Ensemble Learning over the Network
For the benefit of the reader, we briefly report a few basic concepts, referring to the aforementioned article for the details of the hypothesized architecture.
@@
It is envisaged that through ad-hoc functions included in the first level, the ML model to be distributed in the PDP is trained and then properly broken down into individual WLs coded as WL-VNFs that will cooperate with one another to perform attack detection and response. These WL-VNFs are made available to the orchestration functions that are in the second level. Here, an optimal distribution strategy of the WL-VNFs among PDP devices is decided, which allows the selected switches that host the WL-VNFs to operate cooperatively as an active IDS in the network. The whole process is depicted in Fig.1.
The activities described in this paper refer to what was only theorized at the second level of the mentioned architecture, but not previously developed. Specifically, the goal is to find the set of WL-VNFs and the switches that shall host them in such a way as to maximize the security coverage of the considered network, i.e., the effectiveness in detecting and reacting to the maximum number of attacks.
@@
The flow, as it passes through the network, is analyzed by the various WLs that constitute the SL, and each one signals the result of its inference to the others. If a WL realizes that it is the last of the set that constitutes an SL and that all the others have already performed the flow analysis, it completes its analysis, and through a majority voting algorithm takes the final decision, blocking the flows that the WLs chain deems malicious. The algorithm is completely distributed and does not require human involvement or intervention of the network controller. To allow distributed WL-VNFs to inform each other on the inferences carried out for a network flow, a custom header, P4-encoded, is considered as well as a procedure carried out by the PDP device augmented with the WL-VNFs.
IV Problem Formulation
@@
Without losing generality, we will refer to a specific type of vulnerability of the programmable network considered, namely a DDoS attack in which the attacking hosts could also be compromised devices operating within the network. Therefore, it is assumed that the attacks can come not only from outside the network and therefore counteracted with a perimeter defense system (traditional, such as a firewall, or even innovative such as a entire AI model embedded in the entry/exit node of the network) but also from hosts inside the network that is intended to be protected. The specific optimization does not depend on the type of attacks – that is instead related to the source of data used to train the SL and its WLs. The obtained results in terms of efficiency and effectiveness can, therefore, be generalized to other type of attacks provided that the SL is trained on traffic profile patterns that encode the attacks that one desires to detect.
IV-A SL model decomposition into component WLs
As previously mentioned, the proposed split-AI technique resorts to the Ensemble Learning theory. More specifically, the core idea of the proposal lies in the disaggregation of complex Ensemble Strong Learner (SL) models. In the context of ensemble learning, a Strong Learner is a model that achieves high predictive performance by combining the outputs of multiple individual weaker learners, i.e., models that perform only slightly better than random guessing (they achieve an accuracy at least greater than 50%). Formally, let $\mathcal{SL}=\{WL_{1},WL_{2},\dots,WL_{T}\}$ be a set of weak learners where each $WL_{i}:\mathcal{X}\rightarrow\mathcal{Y}$ is a model trained on the input space $\mathcal{X}$ and output space $\mathcal{Y}$. In addition, given =$\mathcal{I}=\{f_{1},f_{2},...,f_{n}\}$ the set of input features on which the $\mathcal{SL}$ has been trained. Each $WL_{i}\in\mathcal{SL}$ is trained considering a subset $I_{i}\subset\mathcal{I}$ of input features, such that:
$WL_{i}\ \text{trained on}\ I_{i}\subset\mathcal{I}\ \forall\ WL_{i}\ \in\mathcal{SL},\text{with}$ (1)
$\bigcap_{i=1}^{T}I_{i}\neq\emptyset$
Given the mentioned formulation, a strong learner $SL:\mathcal{X}\rightarrow\mathcal{Y}$ is defined as an aggregation of the weak learners, such that:
$\mathcal{SL}(x)=\mathcal{A}(WL_{1}(x),WL_{2}(x),\dots,WL_{T}(x))$ (2)
where $\mathcal{A}(\cdot)$ is an aggregation function (e.g., majority voting for classification, weighted average for regression) that combines the independent intermediate inferences computed by each single WL in the SL. The number $T$ of WLs that compose the $\mathcal{SL}$ is required to be odd, to ensuring a majority-based aggregation function can always reach a unique decision. It is worth noticing that the SL is constructed through a bagging-based training process, which ensures that the resulting WLs are mutually independent.
The described AI technique naturally aligns with networking principles, particularly resembling the concept of VNF chaining.
In the proposed Projecting the Ensemble Learning over the Network split-AI technique, the $\mathcal{SL}$ is disaggregated and its weak learners are distributed and deployed onto programmable network elements. Let $\mathcal{P}=\{N_{1},N_{2},\dots,N_{t}\}$ be the ordered set of programmable network nodes traversed by a data flow. Each node $N_{i}$ hosts the weak learner $WL_{i}$ – encoded as VNF – such that:
$N_{i}\mapsto WL_{i}(x_{I_{i}})\quad\text{with}\quad x_{I_{i}}=\pi_{I_{i}}(x)$ (3)
where $\pi_{I_{i}}$ is the projection operator selecting the subset $I_{i}$ of features from $x$.
The inference result is computed incrementally by each involved device in a distributed fashion. The final decision is produced by the last traversed node $N_{t}$ of the path, which applies the aggregation function:
$\mathcal{SL}(x)=A\left(\{WL_{i}(x_{I_{i}})\}_{i=1}^{t}\right)=A\left(r_{1},r_{2},\dots,r_{t}\right)$ (4)
computed at $N_{t}$. This is achieved through a custom packet header that carries the intermediate results computed by each WL-VNF, allowing subsequent components to access these partial outputs and ultimately reconstruct the final result of the original SL. Notably, the WLs composing the SL are mutually independent, and as such, no ordering constraints are required during their traversal. This means that once a node $N_{t}$ has all the required intermediate inferences, it compute the $A(\cdot)$ to aggregate them and produce the final outcome.
IV-B Optimization problem definition
@@
Given the described Split-AI technique, once a complex SL is disaggregated into its lighter individual but functional components, an optimal deployment strategy is needed in order to place them on programmable networking devices accounting for a trade-off between network security and effective network management. In this direction, we represent a network by using a graph. The nodes in our model represent the network nodes in which the WL-VNFs can be deployed, while the edges denote the links between network elements. We use node coloring to represent the implementation of specific WL-VNFs, where each color [^1] The terms “color” and “SL/WL-VNF” will be used interchangeably. More specifically, SL-VNF refers to a scenario in which only one color is needed. corresponds to a different type of WL-VNF and the coloring cost corresponds to the associated implementation cost. For instance, an SL composed of three WLs will determine three WL-VNFs and therefore three different colors (e.g., red, green, and blue), as shown in Fig.2.